OpenStudy (ashking1):

Find a cubic function with the given zeros. 6, -3, 5

OpenStudy (dannyo19):

When you have found the roots or zeros of polynomials they can be written as factors... Like (x+2)(x-3)= x^2-x-6

OpenStudy (ashking1):

im sorry im very lost lol

OpenStudy (dannyo19):

What would be the zeros for (x+2)(x-3)?

OpenStudy (ashking1):

like wha happens if i foil itt

OpenStudy (dannyo19):

They would be -2 and 3 because the zeros are when y=0

OpenStudy (dannyo19):

0=(x+2)(x-3)

OpenStudy (ashking1):

ok i think get it alittle

OpenStudy (ashking1):

so six can be x-6

OpenStudy (dannyo19):

See when I make x=-2 or 3 then y does in fact equal zero

OpenStudy (dannyo19):

y=(x+2)(x-3)

OpenStudy (dannyo19):

So now in order to get a cubic function you'd have to have x times x times x to get x^3

OpenStudy (dannyo19):

Then how about (x+a)(x+b)(x+c) form?

OpenStudy (dannyo19):

Your zeros are 6 , -3 and 5 so for each factor x will equal 6 then -3 then 5 in order to get zero with that factor...

OpenStudy (ashking1):

ok so i think i git it is it x^3-8x^2-3x+90

OpenStudy (dannyo19):

What factors did you use?

OpenStudy (dannyo19):

How do you make 6 equal zero, -3 equal zero and 5 equal zero, get it?

OpenStudy (ashking1):

x-6 x+3 and x-5

OpenStudy (dannyo19):

(x-6)(x+3)(x-5) is a factored form of a cubic function do you need to expand it?

OpenStudy (dannyo19):

If the question doesn't say expand that's fine that way..

OpenStudy (ashking1):

more like i foiled it

OpenStudy (dannyo19):

Ok, if you need to do that you can...

OpenStudy (ashking1):

alright so i foiled it and it gave me x^3-8x^2-3x+90

OpenStudy (dannyo19):

Looks good

OpenStudy (ashking1):

thank you

OpenStudy (dannyo19):

It's also called expand (x-6)(x+3)(x-5) if you want to get answers fast on calculators alternatively to the foil method.